# SIDIS at small at next-to-leading order: Gluon contribution

### v2

@@ I Introduction
Deep inelastic scattering (DIS) is the cleanest environment in which to probe internal partonic structure of hadrons and nuclei [1, 2]. Due to the experimentally observed fast rise of gluons at small Bjorken one expects that a high energy hadron or nucleus is a maximum occupancy state of predominantly gluons [3, 4]. This high density gluonic state is referred to as the color glass condensate (CGC) [5, 6, 7] and is expected to give the dominant contribution to scattering cross sections at high energy. While there have been tantalizing hints of this saturated state at HERA, RHIC and the LHC [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 23, 49, 50, 51, 52, 53, 54] its presence is far from being firmly established (see [55, 56, 57, 58, 59] for reviews). The proposed Electron Ion Collider (EIC) will enable us to look for gluon saturation in various channels with unprecedented precision and accuracy. Among the most promising processes in which gluon saturation is expected to play a dominant role are single and double inclusive jet and hadron production. Measurements of transverse momenta, angular correlations and rapidity dependence of these processes will shed light on dynamics of gluons saturation and its energy dependence and enable a detailed comparison with theoretical expectations based on CGC formalism [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. To improve the theoretical accuracy of CGC-based predictions there has been significant progress made in calculating next to leading order (NLO) and beyond eikonal corrections to observable which are sensitive to gluon saturation [70, 71, 72, 73, 74, 75, 76, 77, 78, Caucal:2022ulg, Caucal:2023nci, Caucal:2023fsf, 79, Fucilla:2023mkl, Fucilla:2022wcg, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, Iancu:2020jch, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, Mukherjee:2023snp, Fu:2023jqv, Boussarie:2021wkn, Boussarie:2023xun].

### v3

@@ I Introduction
Deep inelastic scattering (DIS) is the cleanest environment in which to probe internal partonic structure of hadrons and nuclei [1, 2]. Due to the experimentally observed fast rise of gluons at small Bjorken one expects that a high energy hadron or nucleus is a maximum occupancy state of predominantly gluons [3, 4]. This high density gluonic state is referred to as the color glass condensate (CGC) [5, 6, 7] and is expected to give the dominant contribution to scattering cross sections at high energy. While there have been tantalizing hints of this saturated state at HERA, RHIC and the LHC [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 23, 49, 50, 51, 52, 53, 54] its presence is far from being firmly established (see [55, 56, 57, 58, 59] for reviews). The proposed Electron Ion Collider (EIC) will enable us to look for gluon saturation in various channels with unprecedented precision and accuracy. Among the most promising processes in which gluon saturation is expected to play a dominant role are single and double inclusive jet and hadron production. Measurements of transverse momenta, angular correlations and rapidity dependence of these processes will shed light on dynamics of gluons saturation and its energy dependence and enable a detailed comparison with theoretical expectations based on CGC formalism [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. To improve the theoretical accuracy of CGC-based predictions there has been significant progress made in calculating next to leading order (NLO) and beyond eikonal corrections to observable which are sensitive to gluon saturation [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130].